Find: (3/7) ÷ (8/7) Find: (3/7) ÷ (8/7) Find: (3/7) ÷ (8/7)

Find:
(3/7) ÷ (8/7)
Find:
(3/7) ÷ (8/7)
Find:
(3/7) ÷ (8/7)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We have,
= (3/7) × reciprocal of (8/7)
= (3/7) × (7/8)
= (3 × 7) / (7 × 8)
= (3 × 1) / (1 × 8)
= 3/8

Thanks